# ChartMimic: Evaluating LMM’s Cross-Modal Reasoning Capability via Chart-to-Code Generation

### v1

@@ 2.5 Comparisons with Existing Benchmarks
To further distinguish the difference between ChartMimic and other existing ones, we elaborate the benchmark details in Table 2. From the chart understanding perspective, the prior benchmarks [21, 22, 23] are heavily focused on simple questions about the data in the charts and lack of both difficulty and diversity in chart types. In the code generation aspect, HumanEval [6], MBPP [7] and MatPlotBench [20] only consider tasks with text inputs, which may not meet requirements in the era of LMMs. Recently, MMCode [24] attempted to create a benchmark for multimodal code generation, but the vision inputs for their task are still overly simple and only have a single pass rate evaluation metric. Design2Code [25] is the most similar one to ours, which focuses on automated webpage generation. Although they use multi-level evaluation metrics like ours, their test HTMLs all come from the C4 dataset [27], which may pose a risk of data leakage. Our ChartMimic benchmark gives complex scientific charts as inputs, demanding capabilities on grounding visual understanding into code generation, and provides multi-level evaluation metrics.
@@ 3.1 Baseline Setup
We benchmark $14$ widely utilized proprietary and open-source models currently available in the field. For proprietary models, we consider $3$ representative models: GPT-4V [2], Claude-3-opus [28] and GeminiProVision [3]. For the open-weight models, we choose $11$ competitive models. We list them in increasing order of total parameter size from 4.2B to 34.8B, which are: Phi-3-Vision [29], LLaVA-Next-Vicuna-7B [30], DeepSeek-VL-7B [31], LLaVA-Next-Mistral-7B [30], IDEFICS2-8B [32], MiniCPM-Llama3-V 2.5 [33], Qwen-VL-Chat [34], LLaVA-Next-Vicuna-13B [30], Cogvlm2-llama3-chat-19B [35], InternVL-Chat-V1.5 [36], LLaVA-Next-Yi-34B [30]. We start with the direct prompting method, which provides the reference chart with direct instructions. The specific instructions and models’ decoding setups can be found in Appendix D.
3.2 Main Results
@@
Text Layout Type Color Avg. GPT-4V
@@
GeminiProVision - 64.2 36.9 58.6 45.1 39.6 45.0 38.1 41.6
Claude-3-opus - 86.4 61.8 79.0 45.3 37.9 56.0 45.4 50.7
GPT-4V - 91.4 74.3 87.1 73.8 62.1 74.3 68.4 71.4
@@
Qwen-VL-Chat 9.6B 4.0 0.5 2.8 1.1 0.8 1.3 0.9 1.1
LLaVA-Next-Vicuna-7B 7.1B 6.2 0.8 2.1 1.4 1.6 1.5 1.6 1.6
LLaVA-Next-Vicuna-13B 13.4B 11.6 1.9 5.7 4.0 4.3 4.0 2.8 3.4
IDEFICS2-8B 7.6B 42.4 5.7 28.4 11.9 10.9 14.2 8.8 11.5
LLaVA-Next-Yi-34B 34.8B 45.6 11.4 26.7 18.2 16.6 18.2 12.7 15.4
LLaVA-Next-Mistral-7B 7.6B 57.6 11.3 31.5 22.3 21.9 21.8 14.8 18.3
DeepSeek-VL-7B 7.3B 39.0 13.8 26.4 23.9 19.0 20.8 16.8 18.8
Cogvlm2-llama3-chat-19B 19.2B 51.4 20.6 31.1 21.7 20.0 23.3 17.4 20.4
InternVL-Chat-V1.5 26.0B 60.0 26.9 42.1 26.1 24.9 30.0 23.4 26.7
MiniCPM-Llama3-V2.5 8.4B 58.2 22.5 40.1 31.4 26.6 30.2 24.2 27.2
Phi-3-Vision 4.2B 66.0 36.4 45.0 39.0 34.3 38.7 25.1 31.9
@@
We present the main results of 14 LMMs on ChartMimic. Table 3 and Table 4 show the results on the Direct Mimic and Customized Mimic task, respectively. The key findings are as follows:
@@
GPT-4V performs best among proprietary models, while Phi-3-Vision excels among open-weight models. In proprietary models, GPT-4V achieves an overall score of $71.4$ in Direct Mimic and $72.33$ in Customized Mimic, surpassing the second-place model Claude-3-opus by $21.3$ and $21.1$, respectively. Within open-weight models, Phi-3-Vision, despite its 4.2B parameters, reaches the overall score of $31.9$ in Direct Mimic and $40.18$ in Customized Mimic, which outperforms LLaVA-Next-Yi-34B ($15.4$ in Direct Mimic and $20.2$ in Customized Mimic). This indicates that even models with fewer parameters can achieve decent performance through a refined training process.
@@
A performance gap persists between open-weight LMMs and proprietary ones. Although several open-weight LMMs can exhibit performance comparable to GPT-4V across various benchmarks [12], even the best-performing Phi-3-Vision falls short of achieving half the performance of GPT-4V on both Direct Mimic and Customized Mimic. This apparent performance disparity demonstrates the challenging nature of our ChartMimic benchmark for current open-weight LMMs. Moreover, open-weight LMMs even exhibit notable deficiencies in generating executable code, with a majority of them showing an execution rate below $60\%$. These findings emphasize that there is still a considerable scope for the open-source community to enhance LMMs’ capabilities in terms of complex visual understanding, code generation and cross-modal reasoning.
@@
When given additional user-customized data, LMMs exhibit performance improvements. In the Customized Mimic task, most LMMs get performance improvement compared to the Direct Mimic task, particularly in terms of Text Score. This enhancement can be attributed to the provision of user-customized data, which alleviates the burden on LMMs to recognize textual and data information within the charts. Notably, InternVL-Chat-V1.5 experiences a decrease in its overall score and execution rate. This decline may arise from the reason that user-customized data increases the burden on the models to process different modalities of information. Overall, except for GPT-4V, the performance of the remaining LMMs is still below 60.0 in the Customized Mimic task. This indicates that, in addition to the capability to recognize data in the chart, understanding the layout of the chart and code generation are also key factors for LMMs to better accomplish the task.
@@
Text Layout Type Color Avg. GPT-4V
@@
GeminiProVision - 74.0 47.2 64.9 56.0 46.4 53.6 46.8 50.2
Claude-3-opus - 88.2 65.2 82.8 48.2 39.6 58.9 54.4 56.7
GPT-4V - 93.2 77.4 89.2 78.3 65.6 77.6 72.1 74.9
@@
Qwen-VL-Chat 9.6B 6.8 2.6 3.8 3.5 2.6 3.2 2.2 2.7
LLaVA-Next-Vicuna-7B 7.1B 8.4 2.2 4.8 3.6 3.0 3.4 2.7 3.1
LLaVA-Next-Vicuna-13B 13.4B 32.2 9.3 21.5 17.4 14.2 15.6 10.8 13.2
Cogvlm2-llama3-chat-19B 19.2B 37.8 19.1 25.6 17.6 17.1 19.9 15.0 17.4
LLaVA-Next-Mistral-7B 7.6B 48.0 21.8 30.4 22.0 21.7 24.0 17.4 20.7
IDEFICS2-8B 7.6B 52.4 23.6 33.5 24.1 16.2 24.4 19.8 22.1
LLaVA-Next-Yi-34B 34.8B 59.4 33.4 45.1 26.2 24.9 32.4 17.7 25.0
InternVL-Chat-V1.5 26.0B 48.6 30.1 38.2 27.4 22.5 29.5 23.3 26.4
DeepSeek-VL-7B 7.3B 55.2 28.2 45.8 40.4 33.5 37.0 28.4 32.7
MiniCPM-Llama3-V2.5 8.4B 69.2 40.0 53.9 43.1 35.8 43.2 32.1 37.7
Phi-3-Vision 4.2B 67.2 30.9 52.5 46.6 42.0 43.0 35.7 39.4
@@ 4.1 Different Complexity Levels
Figure 3: Overall scores of top-$6$ performed models at different complexity levels.
As shown in Sec. 2.3, the performance of the top-$6$ models across different complexity levels in ChartMimic is reported. With increasing difficulty, there is a consistent decline in performance across all tasks, resembling a stepped pattern. For example, in Customized Mimic task, the performance of GPT-4V at easy, medium, and hard levels are $82.8$, $69.8$, and $61.6$, respectively. Further observations across all tasks indicate that the performance decline of open-source models from easy to medium difficulty is more pronounced than from medium to hard. In contrast, proprietary models exhibit a consistent rate of decline. This suggests that the foundational capabilities of the open-source models are comparable, and the observed performance variations may largely be influenced by the quality of data. These results demonstrate that the benchmark tasks are inherently challenging and confirms the efficacy of the established difficulty levels.
@@ 4.2 Different Prompting Methods
Table 5: Results for different prompting methods on GPT-4V on Direct Mimic task.
Model Low-Level High-Level Overall
Text Layout Type Color Avg. GPT-4V
Direct 74.3 73.8 87.1 62.1 74.3 68.4 71.4
HintEnhanced 71.0 71.6 85.5 59.1 71.8 67.3 69.6
SelfRevision 71.0 70.8 84.3 59.9 71.5 65.1 68.3
Scaffold 67.4 67.5 81.7 56.3 68.2 64.4 66.3
We further explore the impact of different prompting methods. Specifically, we choose the best-performing model, GPT-4V [2], to study its performance on the Direct Mimic task. We select three representative methods for prompting: HintEnhanced, SelfRevison, and Scaffold Prompting [37]. HintEnhanced uses prompt with chain-of-thought [38], explicitly prompting the model to pay attention to important details (e.g., layout, type, text, etc). SelfRevision involves inputting the model’s own output and the rendered chart as additional information, hoping that the model can self-reflect its results and then produce better. Scaffold Prompting [37] which overlays a dot matrix within the image as visual information anchors and leverages multi-dimensional coordinates as textual positional references. Detailed introductions of these prompting methods can be found in Appendix D.
As demonstrated in Table 5, the Direct method proves to be the most straightforward and effective for GPT-4V. Other methods, to varying degrees, tend to reduce performance. Upon examining the cases, we find that the HintEnhanced method, which involves generating captions first, can introduce hallucinations that lead to errors in the subsequently generated code. Case studies suggest that although SelfRevision can identify discrepancies between output results, it often provides incorrect suggestions, resulting in subsequent errors. As for the Scaffold, its implementation can interfere with existing coordinate axis information in charts with high information density, thereby negatively impacting performance. We provide error cases of three prompting methods in Appendix D.
@@ 4.3 Correlation with Human Evaluation
High-Level 0.6942 $<0.0001$
Low-Level 0.7538 $<0.0001$
To evaluate the reliability of the proposed automatic metrics, we calculate their correlation with human evaluations. Specifically, we collect 1,200 charts collected from GPT-4V using four different prompting methods (Sec. 4.2) on Direct Mimic. Each charts is evaluated by three individual evaluators, who assign scores ranging from 00 to $100$ based on the similarity with the ground-truth chart. Details about the human evaluation process are provided in Appendix E. As shown in Table 6, the Pearson correlation coefficient ($r$) indicates that both the high-level ($r=0.6942$) and low-level metric ($r=0.7538$) have a high correlation with human judgment, demonstrating the reliability of our multi-level metrics. Moreover, the low-level metric demonstrates a higher correlation with human judgments compared to the high-level metric. This is attributed to its dependence on code execution logic, enabling it to capture detailed elements.

### v2

@@ 2.5 Comparisons with Existing Benchmarks
To further distinguish the difference between ChartMimic and other existing ones, we elaborate the benchmark details in Tab. 2. From the chart understanding perspective, the prior benchmarks (Masry et al., 2022; Kantharaj et al., 2022; Rahman et al., 2023; Wang et al., 2024b) are only focused on questions about the data in the charts without necessitating the advanced cognitive capabilities of LMMs, which include visual understanding, code generation, and cross-modal reasoning. In the code generation aspect, HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021) and MatPlotBench (Yang et al., 2024) only consider tasks with text inputs, which may not meet requirements in the era of LMMs. Recently, MMCode (Li et al., 2024b) attempted to create a benchmark for multimodal code generation, but the vision inputs for their task are still overly simple and only have a single pass rate evaluation metric. Design2Code (Si et al., 2024) and Plot2Code (Wu et al., 2024a) are the most similar ones to ours. Although they use multi-level evaluation metrics like ours, their test is crawled directly from the internet or existing datasets, which may pose a risk of data leakage. Our ChartMimic benchmark gives complex scientific charts as inputs, demanding capabilities on grounding visual understanding into code generation, and provides multi-level evaluation metrics.
@@ 3.1 Baseline Setup
We benchmark $17$ widely utilized proprietary and open-source models currently available in the field. For proprietary models, we consider $3$ representative models: GPT-4o (OpenAI, 2024), Claude-3-opus (Anthropic, 2024) and GeminiProVision (Google, 2023). For the open-weight models, we choose $14$ competitive models with total parameter size from 2.2B to 76.0B: InternVL2(2B, 4B, 8B, 26B, 76B) (Chen et al., 2023), Qwen2-VL(2B, 7B) (Wang et al., 2024a), Phi-3-Vision (phi, 2024), DeepSeek-VL-7B (Lu et al., 2024), LLaVA-Next(7B, 34B) (Li et al., 2024a), IDEFICS2-8B (Laurençon et al., 2024), MiniCPM-Llama3-V2.5 (Xu et al., 2024) and Cogvlm2-llama3-chat-19B (Wang et al., 2023a). We start with direct prompting, which provides the reference chart with direct instructions. The specific instructions and model configurations can be found in Appendix F.
@@
Text Layout Type Color Avg. GPT-4o
@@
GeminiProVision - 68.2 52.6 64.2 51.3 47.1 53.8 53.3 53.6
Claude-3-opus - 83.3 66.8 83.1 49.9 42.1 60.5 60.1 60.3
GPT-4o - 93.2 81.5 89.8 77.3 67.2 79.0 83.5 81.2
@@
IDEFICS2-8B 7.6B 49.0 6.2 33.1 9.2 9.0 14.4 17.6 16.0
DeepSeek-VL-7B 7.3B 41.3 15.3 26.6 19.7 14.5 19.0 20.4 19.7
LLaVA-Next-Yi-34B 34.8B 50.2 15.9 29.6 17.6 15.2 19.6 20.6 20.1
LLaVA-Next-Mistral-7B 7.6B 59.7 14.0 31.1 19.8 17.8 20.7 21.3 21.0
Qwen2-VL-2B 2.6B 47.0 20.1 29.5 21.3 17.9 22.2 23.4 22.8
Cogvlm2-llama3-chat-19B 19.2B 50.5 21.3 31.8 18.4 17.0 22.1 24.5 23.3
InternVL2-2B 2.2B 52.5 23.6 35.8 16.0 15.4 22.7 24.2 23.5
Qwen2-VL-7B 8.2B 67.0 26.4 51.0 31.0 23.3 32.9 35.0 34.0
InternVL2-4B 4.2B 66.2 34.7 51.7 25.2 23.6 33.8 38.4 36.1
InternVL2-8B 8.1B 61.8 31.5 51.1 28.6 26.2 34.4 38.9 36.6
MiniCPM-Llama3-V-2.5 8.4B 80.3 30.7 49.6 38.6 27.6 36.6 42.1 39.4
Phi-3-Vision-128K 4.2B 66.7 37.5 49.6 37.4 29.8 38.6 41.0 39.8
InternVL2-26B 26.0B 69.3 39.2 58.7 35.9 31.8 41.4 47.4 44.4
InternVL2-Llama3-76B 76.0B 83.2 54.1 74.5 49.2 41.5 54.8 62.2 58.5
3.2 Main Results
@@
We present the main results of $17$ LMMs on ChartMimic. Tab. 3 and Tab. 4 show the results on the Direct Mimic and Customized Mimic task, respectively. The key findings are as follows:
@@
GPT-4o performs best among proprietary models, while InternVL2-Llama3-76B excels among open-weight models. In proprietary models, GPT-4o achieves an overall score of $81.2$ in Direct Mimic and $83.2$ in Customized Mimic. Within open-weight models, InternVL2-Llama3-76B, reaches the overall score of $58.5$ in Direct Mimic and $64.7$ in Customized Mimic, which have compareble perfromance with proprietary model like Claude-3-opus ($60.3$ in Direct Mimic and $65.4$ in Customized Mimic). Meanwhile, Phi-3-Vision-128K despite its 4.2B parameters, reaches the overall score of $39.8$ in Direct Mimic and $42.1$ in Customized Mimic, which outperforms LLaVA-Next-Yi-34B ($20.1$ in Direct Mimic and $35.3$ in Customized Mimic). This indicates that even models with fewer parameters can achieve decent performance through a refined training process.
@@
A performance gap persists between open-weight LMMs and proprietary ones. Although several open-weight LMMs can exhibit performance comparable to GPT-4o across various benchmarks (OpenCompass, 2023), even the best-performing InternVL2-Llama3-76B falls short of achieving the performance of GPT-4o both Direct Mimic and Customized Mimic. This apparent performance disparity demonstrates the challenging nature of our ChartMimic benchmark for current open-weight LMMs. Moreover, open-weight LMMs even exhibit notable deficiencies in generating executable code, with a majority of them showing an execution rate below $75\%$. These findings highlight that there is still a considerable scope for the open-source community to enhance LMMs’ capabilities regarding complex visual understanding, code generation and cross-modal reasoning.
@@
When given additional user-customized data, LMMs exhibit performance improvements. In the Customized Mimic task, most LMMs get performance improvement compared to the Direct Mimic task, particularly in terms of Text Score. This enhancement can be attributed to the provision of user-customized data, which alleviates the burden on LMMs to recognize textual and data information within the charts. Notably, Cogvlm2-llama3-chat-19B and Qwen2-VL-2B experienced a decrease in their overall score and execution rate. This decline may arise from the reason that user-customized data increases the burden on the models to process different modalities of information. Overall, except for GPT-4o, the performance of the remaining LMMs is still below $65.0$ in the Customized Mimic task. This indicates that, in addition to the capability to recognize data in the chart, understanding the layout of the chart and code generation are also key factors for LMMs to better accomplish the task.
@@
Text Layout Type Color Avg. GPT-4o
@@
GeminiProVision - 76.2 52.2 70.9 56.0 49.4 57.1 59.6 58.4
Claude-3-opus - 88.2 75.2 86.8 54.1 44.3 65.1 65.7 65.4
GPT-4o - 96.5 88.5 92.9 79.2 67.6 82.1 84.3 83.2
@@
Qwen2-VL-2B 2.6B 35.8 17.4 23.9 19.7 16.5 19.4 21.4 20.4
Cogvlm2-llama3-chat-19B 19.2B 38.7 19.0 27.9 16.5 15.7 19.8 21.6 20.7
LLaVA-Next-Mistral-7B 7.6B 49.0 20.0 32.0 22.6 19.9 23.6 24.7 24.2
IDEFICS2-8B 7.6B 49.2 21.6 32.2 18.1 12.2 21.0 27.3 24.2
InternVL2-2B 2.2B 49.3 22.2 35.4 20.0 18.1 23.9 27.8 25.9
LLaVA-Next-Yi-34B 34.8B 64.2 28.7 44.8 32.9 27.7 33.5 37.1 35.3
DeepSeek-VL-7B 7.3B 59.3 27.5 47.5 36.8 31.5 35.8 39.3 37.6
Phi-3-Vision-128K 4.2B 67.8 29.7 52.5 42.3 36.5 40.3 44.0 42.1
InternVL2-4B 4.2B 74.0 41.3 55.6 39.6 33.1 42.4 47.8 45.1
Qwen2-VL-7B 8.2B 73.3 41.0 56.3 43.5 34.2 43.8 47.8 45.8
InternVL2-8B 8.1B 73.0 43.1 54.4 39.9 35.4 43.2 48.9 46.1
MiniCPM-Llama3-V-2.5 8.4B 78.7 40.8 58.0 44.8 33.2 44.2 51.5 47.9
InternVL2-26B 26.0B 73.7 43.9 62.3 43.5 34.3 46.0 51.1 48.6
InternVL2-Llama3-76B 76.0B 89.8 57.8 79.0 63.5 50.5 62.7 66.7 64.7
@@ 4.1 Different Complexity Levels
Figure 3: Overall scores of models at different complexity levels.
We report the performance of the top-$6$ models across different complexity levels for ChartMimic in  Fig. 3. There is a consistent decline in performance across all tasks as the difficulty increases. For example, in Direct Mimic task, the performance of GPT-4o at easy, medium, and hard levels are $86.5$, $77.7$, and $74.8$, respectively. These results demonstrate that ChartMimic are inherently challenging and confirm the efficacy of the established difficulty levels. Additionally, for weak LMMs such as MiniCPM-Llama3-V-2.5, providing customized data at the Easy Level can help it perform tasks better (from $46.8$ to $59.7$). However, at the Hard Level, even when given customized data, their performance do not show improvement (from $31.9$ to $31.5$). This indicates that for Easy Level, complex reasoning abilities are not required, and providing data is sufficient to complete the task well. However, for Hard Level charts, LMMs need complex code generation capabilities grounded on visual understanding. In this case, even when given data, if the LMMs cannot perform the corresponding code reasoning, they still cannot complete the task. Moreover, when given addtional data, if the data is too long and complex, the LLM may still become confused, leading to a certain decrease in performance.
@@ 4.2 Different Prompting Methods
Table 5: Results for different prompting methods for GPT-4o and different size of InternVL2 LMMs on Direct Mimic task. The best scores for each model are in bold.
Model Params Method Exec. Rate Low-Level High-Level Overall
Text Layout Type Color Avg. GPT-4o
GPT-4o - Direct 93.2 81.5 89.8 77.3 67.2 79.0 83.5 81.2
HintEnhanced 92.0 80.9 89.7 78.2 67.9 79.2 83.0 81.1
SelfReflection 95.2 86.6 93.7 81.1 68.2 82.4 87.1 84.8
Scaffold 91.0 81.3 90.6 75.1 63.9 77.7 80.3 79.0
InternVL2-Llama3-76B 76.0B Direct 83.2 54.1 74.5 49.2 41.5 54.8 62.2 58.5
HintEnhanced 78.0 54.3 71.7 50.2 41.6 54.5 60.3 57.4
SelfReflection 82.5 57.0 76.2 56.4 45.0 58.7 63.6 61.1
Scaffold 76.7 52.1 71.4 49.1 40.5 53.3 60.1 56.7
InternVL2-26B 26.0B Direct 69.3 39.2 58.7 35.9 31.8 41.4 47.4 44.4
HintEnhanced 66.8 40.4 54.8 36.6 33.0 41.2 47.0 44.1
Self 69.0 39.5 57.5 35.6 30.5 40.8 47.1 43.9
Scaffold 64.3 36.0 53.5 35.4 30.3 38.8 44.0 41.4
InternVL2-8B 8.1B Direct 61.8 31.5 51.1 28.6 26.2 34.4 38.9 36.6
HintEnhanced 58.7 30.1 38.3 30.1 27.5 31.5 35.9 33.7
SelfReflection 56.8 27.4 39.8 25.1 22.9 28.8 33.4 31.1
Scaffold 61.8 22.8 43.5 24.8 20.2 27.8 32.0 29.9
InternVL2-2B 2.1B Direct 52.5 23.6 35.8 16.0 15.4 22.7 24.2 23.5
HintEnhanced 41.5 17.3 19.4 14.3 12.9 16.0 18.5 17.2
SelfReflection 35.5 16.3 22.2 12.8 14.8 16.5 17.3 16.9
Scaffold 38.0 6.5 17.8 3.2 2.2 7.4 15.6 11.5
We further examine the impact of different prompting methods on the performance of ChartMimic benchmark. Specifically, we choose the GPT-4o and InternVL2 series LMMs (2B, 8B, 26B, 76B) to study their performance on the Direct Mimic task. We select three representative prompting methods: HintEnhanced, SelfReflection, and Scaffold Prompting. HintEnhanced uses prompt with chain-of-thought (Wei et al., 2022), explicitly prompting the LMMs to pay attention to important details (e.g., layout, type, text, etc). SelfReflection (Shinn et al., 2024) involves inputting the LMMs’ own output and the corresponding rendered chart as additional information, instructing the LMM’ to self-reflect their output. Scaffold Prompting (Lei et al., 2024) overlays a dot matrix within the figure as visual information anchors and leverages multi-dimensional coordinates as textual positional references. We detail the experimental setups in Appendix F.
As shown in Tab. 5, SelfReflection enables GPT-4o and InternVL2-Llama3-76B to reflect on and correct their outputs, demonstrating notable improvements over Direct Prompting. This underscore self-reflection as a key manifestation of System 2 reasoning (Sloman, 1996; Kumar et al., 2024). For GPT-4o, SelfReflection enhances Text, Layout, and Type Scores, though Color Score remains unchanged due to persistent challenges in fine color discrimination. However, LMMs with less developed reasoning capabilities (InternVL2-26B, 8B, and 2B) show no improvement or even decline with self-reflection. This indicates that only LMMs with substantial reasoning capabilities can effectively engage in self-reflection and result optimization, underscoring the critical role of System 2 reasoning in addressing ChartMimic’s challenges. Regarding HintEnhanced and Scaffold methods, they do not enhance the performance of models and reduce performance to varying degrees. Upon examining the cases, we find that the HintEnhanced method, which involves generating captions first, can introduce hallucinations that lead to errors in the subsequently generated code. As for the Scaffold method, introduction of an additional dot matrix can interfere with existing coordinate axis information in charts with high information density, thereby negatively impacting performance. These negative effects intensify as LMMs’ reasoning capabilities decrease, further highlighting the importance of advanced reasoning capabilities in handling complex prompting strategies. We provide case studies of three prompting methods in Appendix F.
@@ 4.3 Correlation with Human Evaluation
High-Level 0.7041 $<0.0001$
Low-Level 0.7681 $<0.0001$
To evaluate the reliability of the proposed multi-level metrics, we calculate their correlation with human evaluations. Specifically, we collect $1,200$ charts collected from GPT-4o using four different prompting methods (Sec. 4.2) on Direct Mimic. Each charts is evaluated by three individual evaluators, who assign scores ranging from 00 to $100$ based on the similarity with the ground-truth chart. Details about the human evaluation process are provided in Appendix G. As shown in Tab. 6, the Pearson correlation coefficient ($r$) indicates that both the high-level ($r=0.7041$) and low-level metric ($r=0.7681$) have a high correlation with human judgment, demonstrating the reliability of our multi-level metrics. Moreover, the low-level metric demonstrates a higher correlation with human judgments compared to the high-level metric. This is attributed to its dependence on code execution logic, enabling it to capture detailed elements.